OpenStudy (anonymous):

anyone want to help? it might take a little explaining for me to understand the question.. The diameter of a basketball rim is 18 inches. A standard basketball has a circumference of 30 inches. About how much room is there between the ball and the rim in a shot in which the ball goes in exactly in the center of the rim? Show your work.

OpenStudy (anonymous):

any chance you can draw me up one of those pictures with the rim and then a ball in the middle so i can see it better?

OpenStudy (campbell_st):

do you know how to find the circumference...?

OpenStudy (anonymous):

c= 2 pie R

OpenStudy (anonymous):

but i need the diameter of the ball not the circumfrence right?

OpenStudy (campbell_st):

ok... thats good... so lefts find the diameter of the basketball \[30 = 2 \times \pi \times r\] does that make sense

OpenStudy (anonymous):

9.55?

OpenStudy (campbell_st):

the diagram is